OpenStudy (anonymous):

An initial population of 295 quail increases at an annual rate of 7%. Write an exponential function to model the quail population. What will the approximate population be after 2 years? (1 point) A. f(x) = (295 x 0.07)^x;426 B. f(x) = 295(7)^x; 14,455 C. f(x) = 295(0.07)^x; 338 D. f(x) = 295(1.07)^x; 338 please help me!!

OpenStudy (anonymous):

note that \(7\%=.07\) and to increase a number by \(7\%\) you multiply by \(1.07\)

OpenStudy (anonymous):

to increase a number by \(7\%\) \(x\) times multiply by \((1.07)^x\)

OpenStudy (anonymous):

:/ im still confused

OpenStudy (anonymous):

ok suppose you want to increase a number by 7%, like say you want a final price after a 7% sales tax and lets say the thing you are buying cost $50 then the total price would be \(50+.07\times 50=50+3.5=53.5\) another way to calculate this is by the distributive law \(50+.07\times 50=50(1+.07)=50(1.07)\)

OpenStudy (anonymous):

in other words, to increase a number by 7% you can either a) find 7% of that number and add it to the original number, or b) multiply that number by \(1.07\)

OpenStudy (anonymous):

now if you want to increase a number \(P\) by 7% three times i a row, you can write in one step \[P(1.07)^3\]